OpenStudy (anonymous):

Prove tan(θ/2)=sinθ/(1+cosθ)for θ inq quadrant 1 by filling in the reasons below

OpenStudy (anonymous):



OpenStudy (anonymous):

This may help: http://openstudy.com/study#/updates/4f78f650e4b0ddcbb89e7fe4

OpenStudy (anonymous):

@Rohith15 do you know? :)

OpenStudy (anonymous):

Hey. Just see that post. In the 7th comment by Mertsj. Do you understand that??

OpenStudy (anonymous):

A little bot, but I don't know how to put that into my table. Could you?

OpenStudy (anonymous):

i can but it will take time..!!

OpenStudy (anonymous):

That's ok. I just need to fill in the table. 3 calculations with 3 reasons. I will give you a medal if you help me well!!

OpenStudy (anonymous):

@RadEn hey man. This is the same type of question lol still kind of confused on what to write in the table

OpenStudy (raden):

tan(θ/2) = sin(θ/2)/cos(θ/2) --> reason = identity (tan=sin/cos) sin(θ/2) = sqrt(1-cos(θ))/2 and cos(θ/2) = sqrt(1+cos(θ))/2 ---> reason = half angle formula sqrt(1-cos(θ))/ sqrt(1+cos(θ))/2 = sin(θ)/(1+cos(θ)) --> reason = rationalizing the numerator actually, i hate the table. lol

OpenStudy (anonymous):

lol why do you hate the table ??? :)

OpenStudy (anonymous):

BTW thanks for this!